Suppose the position of an object is given by r = (3.0t2 – 6.0t3 j) m. [DIFFERENTIATION] 

  1. Determine its velocity, v, and acceleration, a, as a function of time. [ANS v = (6.0t i – 18.0t2 j) m/s, a = (6.0 i – 36.0t j) m/s2] 
  2. Determine r and v at time t = 2.5 s. [ANS = (19i – 94j) m, v = (15i – 110j) m/s]  
1
Expert's answer
2021-11-07T19:30:02-0500

1)

"v=\dfrac{dr}{dt}=\dfrac{d}{dt}[(3.0t^2\ i-6.0t^3\ j)\ m]=(6.0t\ i-18.0t^2\ j)\ \dfrac{m}{s},""a=\dfrac{dv}{dt}=\dfrac{d}{dt}[(6.0t\ i-18.0t^2\ j)\ m]=(6.0\ i-36.0t\ j)\ \dfrac{m}{s^2}."

2)

"r(t=2.5\ s)=(3.0\times(2.5\ s)^2\ i-6.0\times(2.5\ s)^3\ j)\ m,""r(t=2.5\ s)=(19.0i-94j)\ m.""v(t=2.5\ s)=(6.0\times2.5\ s\ i-18.0\times(2.5\ s)^2\ j)\ \dfrac{m}{s},""v(t=2.5\ s)=(15i-112j)\ \dfrac{m}{s}."